OpenStudy (anonymous):

radical 3 - radical 6 over radical 3 - radical 6

OpenStudy (anonymous):

please help

OpenStudy (jack1):

radical is the same as square root...? ie radical 3 = \[\sqrt 3\]?

OpenStudy (jack1):

\[\huge \frac {\sqrt 3 - \sqrt 6}{ \sqrt 3 - \sqrt 6}\] is this your equation? @345611math22 ?

OpenStudy (jack1):

if so = 1/1 which equals 1 as any number over itself = 1 ie 100/100 = 1 4/4 = 1 sqrt 2/ sqrt 2 = 1